Given the project network and baseline information below, complete the form to develop a status report for the project at the end of period 4 and the end of period 8 LEGEND ES | ID | EF SL LS DUR LF SL 8 D 12 0 8412 0 0 0 2 6 8 12 F15 0 12 315 0 0 0 7 E 10 2 93 12 2 4 5 9
Time Period Task LF DUR ES SL Budget (PV) 1 23 4 5 67 89 10 11 1213 14 15 400 200 200 2300 1510 1560 900 600 00 500 200 600 200 600 200 400 500 110 300 4 12 12 15 0 360 400 400 400 300 400 200 12 200 100 300 Period PV total 200 200400 900 700 710 500 900 760 600 400400 200 100 300 Cumulative PV total 200 400 800 1,700 2,400 3,110 3,610 4,510 5,270 5,870 6,270 6,670 6,870 6,970 7,270 (Do not round intermediate calculations. Round your table answer entries to the nearest whole number.) End of Period 4 Task CV SV Actual %Complete Finished 50 30 EV PV AC 300 900 500 Cumulative Totals
End of Period 8 Task SV Actual % Complete Finished Finished Finished 25 30 EV AC CV PV 300 2,100 1,600 400 500 Cumulative Totals
Answer #1
1 B C D E F G H
2 End of Period 4
3 Task Actual % Complete EV AC PV CV SV
4 A Finished 400 300 400 100 0
5 B 50 1150 900 700 250 450
6 C 30 453 500 600 -47 -147
7 D 0 0 0 0 0 0
8 E 0 0 0 0 0 0
9 Cumulative Totals 2003 1700 1700 303 303
10
11 End of Period 8
12 Task Actual % Complete EV AC PV CV SV
13 A Finished 400 300 400 100 0
14 B Finished 2300 2100 2300 200 0
15 C Finished 1510 1600 1510 -90 0
16 D 25 390 400 0 -10 390
17 E 30 270 500 300 -230 -30
18 F 0 0 0 0 0 0
19 Cumulative Totals 4870 4900 4510 -30 360

Formula

1 B C D E F G H
2 End of Period 4
3 Task Actual % Complete EV AC PV CV SV
4 A Finished 400 300 400 =D4-E4 =D4-F4
5 B 50 =(200+500+200+600+200+600)*C5/100 900 700 =D5-E5 =D5-F5
6 C 30 =(200+400+500+110+300)*C6/100 500 600 =D6-E6 =D6-F6
7 D 0 =F7*C7/100 0 0 =D7-E7 =D7-F7
8 E 0 =F8*C8/100 0 0 =D8-E8 =D8-F8
9 Cumulative Totals =SUM(D4:D8) =SUM(E4:E8) =SUM(F4:F8) =D9-E9 =D9-F9
10
11 End of Period 8
12 Task Actual % Complete EV AC PV CV SV
13 A Finished 400 300 400 =D13-E13 =D13-F13
14 B Finished 2300 2100 =(200+500+200+600+200+600) =D14-E14 =D14-F14
15 C Finished 1510 1600 =(200+400+500+110+300) =D15-E15 =D15-F15
16 D 25 =(360+400+400+400)*25/100 400 0 =D16-E16 =D16-F16
17 E 30 =(300+400+200)*30/100 500 300 =D17-E17 =D17-F17
18 F 0 0 0 0 =D18-E18 =D18-F18
19 Cumulative Totals =SUM(D13:D18) =SUM(E13:E18) =SUM(F13:F18) =D19-E19 =D19-F19